The Balance Sheet and the Income Statement for ABC Manufacturing Corporation are as follows:

DATA (All amounts in SAR unless otherwise indicated, all sales are on credit and no hard keys.)

Balance Sheet:

ABC Corporation Balance Sheet

2019

Cash

310

Accounts receivable

685

Inventories

600

Other current assets

105

Total current assets

1,700

Gross fixed assets

3,750

Accumulated depreciation

-1,500

Net fixed assets

2,250

Total assets

3,950

LIABILITIES (DEBT) AND EQUITY

Accounts payable

350

Short-term notes payable

250

Total current liabilities

600

Long-term debt

950

Total liabilities

1,550

Common stock (par & paid in capital)

1,050

Retained earnings

1,475

Total common equity

2,400

Total liabilities and equity

3,950

ABC Corporation Income Statement

2019

Sales

6,000

Cost of goods sold

-3,700

Gross profits

2,300

Operating expenses:

Selling and G&A expenses

-820

Depreciation expenses

-340

Total operating expenses

-1,160

Operating income (earnings before interest and taxes)

1,140

Interest expense

-200

Earnings before taxes

940

Income taxes

-230

Net Income

710

Number of shares (thousands)

160

Dividends (thousands)

120

Market price per share

40

Calculate the following ratios:

Operating return on assets

Operating profit margin

Total asset turnover

Fixed asset turnover

Debt ratio

Times interest earned

Return on equity

Earnings per share

Dividends per share

Price/Earnings

Book value per share

Market/Book
Answer #1
Operating return on assets = Operating income / Total assets
Operating return on assets = 1140/3950
Operating return on assets = 28.86%
Operating Profit margin = Operating profit / Net Sales
Operating Profit margin = 1140/6000
Operating Profit margin = 19.00%
Total Assets Turnover = Net Sales / Total Assets
Total Assets Turnover = 6000/3950
Total Assets Turnover = 1.52
Fixed Assets Turnover = Net Sales / Fixed Assets
Fixed Assets Turnover = 6000/2250
Fixed Assets Turnover = 2.67
Debt Ratio = Total Debt / Total Assets
Debt Ratio = 1550/3950
Debt Ratio = 0.39
Times Interest earned = EBIT/ Interest
Times Interest earned = 1140/200
Times Interest earned = 5.7
Return on equity = Net income / Shareholders equity
Return on equity = Net income / (total assets - total liabilities)
Return on equity = 710/(3950-1550)
Return on equity = 29.58%
Earning per share = Net income / nos of shares
Earning per share = 710000/160000
Earning per share = 4.44 SAR
Dividends per share = Dividends / Nos of shares
Dividends per share = 120000/160000
Dividends per share = 0.75 SAR
Price to earning = Price per share / Earning per share
Price to earning = 40/4.44
Price to earning = 9.01
Book value per share = Shareholders equity / nos of shares
Book value per share = (total assets - total liabilities)/ nos of shares
Book value per share = (3950-1550)/160
Book value per share = 15
Market to book = Market price per share / Book value per share
Market to book = 40/15
Market to book = 2.67